# On Mirzakhani’s twist torus conjecture

### v2

@@ 2.2. The earthquake flow
By the uniformization theorem, $\mathcal{T}_{g}$ can also be identified with the space of marked hyperbolic structures on $S$. This viewpoint yields global analytic Fenchel–Nielsen coordinates on $\mathcal{T}_{g}$ adapted to any pair of pants decomposition $\gamma$, comprised of $33g-$ 3 coordinate length functions $\ell_{i}$ and $33g-$ 3 real valued twist functions $\tau_{i}$. A choice of Fenchel–Nielsen coordinates requires a choice of what “zero twist” should mean, however, the differentials $d\tau_{i}$ are defined. In other words, the twisting data form a principal $\mathbb{R}^{3g-3}$-bundle over the length data.
@@
There is a flat bundle $\mathcal{P}^{1}\mathcal{T}_{g}$ over $\mathcal{T}_{g}$ with fiber $\mathcal{ML}^{1}(X)$ over $X$; the quotient by $\operatorname{Mod}_{g}$ is denoted by $\mathcal{P}^{1}\mathcal{M}_{g}$. By [Thu98, Theorem 5.1], these bundles can also be identified with the unit cotangent bundles of $\mathcal{T}_{g}$ and $\mathcal{M}_{g}$, respectively. Integrating the fiberwise measures $\hat{\mu}_{\operatorname{Th}}^{X}$ against the Weil–Petersson volume yields the $\operatorname{Mod}_{g}$-invariant, Lebesgue-class Mirzakhani measure on $\mathcal{P}^{1}\mathcal{T}_{g}$. Its local pushforward to $\mathcal{P}^{1}\mathcal{M}_{g}$ is a finite measure $\mu_{\operatorname{Mirz}}$ that we also call by the same name. The total total mass of $\mu_{\operatorname{Mirz}}$ is
@@ 2.3. The conjugacy
Because their leaves are always orthogonal to $\lambda$, the orthogeodesic foliations on the components of $X\setminus\lambda$ can be extended continuously across $\lambda$, yielding a singular foliation $\mathcal{O}_{\lambda}(X)$ on $S$. This foliation also carries a natural transverse measure, defined by assigning to any geodesic arc in $\lambda$ its Lebesgue measure and then extending uniquely to all arcs via transverse isotopy. See [CF24b, Section 5] for more details.
@@
The fact that the length of $\lambda$ on $X$ is equal to the area of $\mathcal{O}(X,\lambda)$ combined with the $\operatorname{Mod}_{g}$-equivariance of $\mathcal{O}$ imply that it induces a bijection
@@
of the same name. The following omnibus theorem combines the main theorems of [Mir08] and [CF24b] together with a special case of the main result of [CF24a].
@@
    $\mathcal{O}$ takes earthquake flow to horocycle flow in a time-preserving way: $\mathcal{O}\circ\operatorname{Eq}_{s}=u_{s}\circ\mathcal{O}$ [^4] Theorem A of [CF24b]; see also Theorem 1.1 of [Mir08].
@@
    Each $k$-pronged singularity of $\operatorname{\mathcal{O}_{\lambda}}(X)$ corresponds to an order $(k-2)$ zero of $\mathcal{O}(X,\lambda)$ and vice versa. [^6] Proposition 5.10 of [CF24b].
@@
    If $\mathcal{Q}$ is a component of a stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$ and $q\in\mathcal{Q}$ has no horizontal saddle connections, then $q$ is a point of continuity for $\mathcal{O}^{-1}|_{\mathcal{Q}}$. [^7] A stronger version of this is proved as Theorem 15.6 of [CF24a].
@@ 3.1. Twist tori
Let $\gamma=(\gamma_{1},\ldots,\gamma_{n})$ be a multicurve on $S$; throughout the paper we use $S\setminus\gamma$ to denote the (possibly disconnected) closed surface obtained by removing an open annulus about each $\gamma_{i}$. Associated to $\gamma$ there is a cut-and-glue fibration that expresses a hyperbolic structure on $S$ in terms of the induced hyperbolic structure on $S\setminus\gamma$ and how these structures are glued together. Let $\mathcal{T}(S\setminus\gamma)$ denote the space of marked [^8] Here, the marking does not record twisting about the boundary. hyperbolic structures with totally geodesic boundary on $S\setminus\gamma$ so that for each $i$, the two boundaries of $S\setminus\gamma$ that are glued along $\gamma_{i}$ have the same length. Then we have the following principal $\mathbb{R}^{n}$-bundle:
@@ 3.2. Arc systems and ribbon graphs
The spine map takes surfaces with boundary lengths $\bm{\ell}$ to ribbon graphs whose boundary lengths are also $\bm{\ell}$. Given a multicurve $\gamma$ on a closed surface $S$, we can therefore piece together the spine maps on each component of $S\setminus\gamma$ to get a homeomorphism
@@
where $\mathcal{TRG}(S\setminus\gamma)$ is defined analogously to $\mathcal{T}(S\setminus\gamma)$; compare [AHC22, pg. 39]. We observe that when $\gamma$ is a pants decomposition, $\operatorname{\mathsf{Sp}}(Y)$ is completely determined by the boundary lengths $\bm{\ell}$. In the general case, the lengths of the $\gamma_{i}$ are explicitly obtainable from the metric on $\operatorname{\mathsf{Sp}}(Y)$, and we will continue to use $\bm{\ell}\in\mathbb{R}^{n}$ to denote the vector of lengths of the curves of $\gamma$.
@@
The space $\mathcal{TRG}(S\setminus\gamma)$ naturally has a piecewise polyhedral structure, in which top-dimensional faces correspond to trivalent metric ribbon graphs and lower-dimensional ones correspond to ribbon graphs with higher valence vertices (this combinatorial structure is also readily apparent in the dual picture in terms of weighted filling arc systems). For a given ribbon graph $\Gamma$ of the same topological type as $S\setminus\gamma$, we set $F(\Gamma)$ to denote the facet of $\mathcal{TRG}(S\setminus\gamma)$ consisting of metrics on $\Gamma$ satisfying the gluing conditions. Luo’s work [Luo07] implies that $\operatorname{\mathsf{Sp}}$ restricts to a real-analytic homeomorphism between each facet and a corresponding subvariety of $\mathcal{T}(S\setminus\gamma)$.
@@ 3.3. Horospheres and slices
Each horosphere has a natural lift $\mathcal{P}H_{\gamma}(\mathbf{h})$ to $\mathcal{P}^{1}\mathcal{T}_{g}$ with $\mathbf{h}\gamma$ as the second coordinate. Both the horosphere and its lift are homeomorphic to $\mathbb{R}^{6g-7}$ and are fibered by twist tori.
Following Forni [For21], we define a horospherical measure to be any probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ that is absolutely continuous with respect to the Lebesgue class with continuous density and such that its restriction to each earthquake flow line is the restriction of Lebesgue to an interval. We also say that the pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$ is horospherical.111111 For example, in [AH21, Liu20, AHC22], such measures are obtained by taking certain $\operatorname{Stab}(\mathbf{h}\gamma)$-invariant measures on $\mathcal{P}H_{\gamma}(\mathbf{h})$, taking a local pushforward to the intermediate quotient $\mathcal{P}^{1}\mathcal{T}_{g}/\operatorname{Stab}(\mathbf{h}\gamma)$, then pushing down to $\mathcal{P}^{1}\mathcal{M}_{g}$.
@@
As before, there is a natural lift to $\mathcal{P}^{1}\mathcal{T}_{g}$, both the horosphere slice and its lift are homeomorphic to a Euclidean space, and both are fibered by twist tori. We say that a probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ is a horosphere slice measure if it is in the class of Lebesgue with continuous density (now on the slice) and the conditional measures on earthquake flow lines are Lebesgue on intervals. Just as above, we also use this term to denote pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$.
@@ 4.1. Twist tori and horospheres
Corresponding to $\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is a flat twist torus parameterizing quadratic differentials built from horizontal cylinders with core curves $\gamma$, heights $\mathbf{h}$, and lengths (circumferences) $\bm{\ell}$; these cylinders are glued with arbitrary twists along the metric ribbon graph $\mathsf{Y}$. Indeed, define
@@
The differentials $q\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ are completely horizontally periodic, because their horizontal foliations are measure equivalent to $\mathbf{h}\gamma$. With definitions as in [Wri15] or §5.1 below, the flat twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is just the orbit of any such by the abelian group $\mathbb{R}^{\gamma}$ of twists in the horizontal cylinders, and $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$ is Lebesgue on $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. We observe that $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is locally an affine submanifold in period coordinates, cut out by equations of the form $\operatorname{Im}(z)=h_{i}$ and $z=\ell(e)$, where $e$ is an edge of $\mathsf{Y}$ with length $\ell(e)$. By our convention that $\bm{\ell}\cdot\mathbf{h}=1$, the intersection of these affine subspaces lies in the unit area locus.
which is actually a lift of the $1$-level set of the extremal length function $L_{\mathbf{h}\gamma}:\mathcal{T}_{g}\to\mathbb{R}_{>0}$ [CF24b, GM91]. Using Theorem 2.1, $\mathcal{Q}H_{\gamma}(\mathbf{h})$ is a leaf of the strong unstable foliation. For any facet of $\mathcal{TRG}(S\setminus\gamma)$ we can also define the extremal length horosphere slice
By construction of $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ and Theorems 2.1 and 2.3, $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ is actually a leaf of the strong unstable foliation in the ambient stratum containing it. The angle around each singularity of the flat metric is $\pi$ times the valence of the corresponding vertex of $\Gamma$; see §2 of [CF24b].
@@
We observe how twist tori and their associated measures interact with the $P$ action.
@@ 4.2. Affine invariant subvarieties
The unit area locus $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$ of a $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant immersed sub-orbifold $\mathbb{R}\mathcal{L}\subset\mathcal{QM}_{g}$ that is locally cut out by $\mathbb{C}$-linear equations defined over $\mathbb{R}$ in period coordinates is called an affine invariant subvariety or AIS. An AIS $\mathcal{L}$ comes equipped with an $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant probability affine measure $\nu_{\mathcal{L}}$, where if $m$ is Lebesgue measure on $\mathbb{R}$, then $dm\,d\nu_{\mathcal{L}}$ is locally an affine linear measure on a finite union of complex subspaces in period coordinates; see [EM18, Definition 1.1]. While we will not use it directly, we mention that Filip proved that every AIS $\mathcal{L}$ is an algebraic variety [Fil16].
@@
Let $\mathcal{Q}$ be any component of any stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$.
@@
    (Measure classification): Every $P$-invariant ergodic measure on $\mathcal{Q}$ is also $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant and is the affine measure on an AIS, of which there are at most countably many.
@@
    (Orbit closures): For any $q\in\mathcal{Q}$, the orbit closure $\overline{P\cdot q}=\overline{\operatorname{\mathsf{PSL}}_{2}\mathbb{R}\cdot q}$ is an AIS.
@@
    (Genericity): For any $q\in\mathcal{Q}$, the orbit $P\cdot q$ equidistributes in its closure $\mathcal{L}$. That is, for any $f\in C_{c}(\mathcal{Q})$ and any $r\in\mathbb{R}\setminus\{0\}$,
[…]
Applying Theorem 4.2 to the setting of Lemma 4.1, we deduce that the time-averaged pushforwards of a twist torus equidistribute to its $P$–orbit closure.
@@
[…]
as $T\to\infty$, where $\nu_{\mathcal{L}}$ is the unique affine probability measure supported on $\mathcal{L}:=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})}.$
@@
$\overline{P\cdot q}=\overline{P\cdot\mathcal{Q}\mathbb{T}^{0}}=\mathcal{L}$

### v3

@@ 2.2. The earthquake flow
By the uniformization theorem, $\mathcal{T}_{g}$ can also be identified with the space of marked hyperbolic structures on $S$. This viewpoint yields global analytic Fenchel–Nielsen coordinates on $\mathcal{T}_{g}$ adapted to any pair of pants decomposition $\gamma$, comprised of 3g33g-3 coordinate length functions $\ell_{i}$ and 3g33g-3 real valued twist functions $\tau_{i}$. A choice of Fenchel–Nielsen coordinates requires a choice of what “zero twist” should mean, however, the differentials $d\tau_{i}$ are defined. In other words, the twisting data form a principal $\mathbb{R}^{3g-3}$-bundle over the length data.
@@
There is a flat bundle $\mathcal{PT}_{g}$ over $\mathcal{T}_{g}$ with fiber $\mathcal{ML}(X)$ over $X$; the quotient by $\operatorname{Mod}_{g}$ is denoted by $\mathcal{PM}_{g}$. Similarly to quadratic differentials, we will decorate with a superscript $1$ when we restrict to the locus of unit-length measured laminations. By [Thu98, Theorem 5.1], these bundles can also be identified with the (unit) cotangent bundles of $\mathcal{T}_{g}$ and $\mathcal{M}_{g}$, respectively. Integrating the fiberwise measures $\hat{\mu}_{\operatorname{Th}}^{X}$ against the Weil–Petersson volume yields the $\operatorname{Mod}_{g}$-invariant, Lebesgue-class Mirzakhani measure on $\mathcal{P}^{1}\mathcal{T}_{g}$. Its local pushforward to $\mathcal{P}^{1}\mathcal{M}_{g}$ is a finite measure $\mu_{\operatorname{Mirz}}$ that we also call by the same name. The total total mass of $\mu_{\operatorname{Mirz}}$ is
@@ 2.3. The conjugacy
Because their leaves are always orthogonal to $\lambda$ and the line field tangent to $\lambda$ is Lipschitz continuous, the orthogeodesic foliations on the components of $X\setminus\lambda$ can be extended continuously across $\lambda$, yielding a singular foliation $\mathcal{O}_{\lambda}(X)$ on $S$. This foliation also carries a natural transverse measure, defined by assigning to any geodesic arc in $\lambda$ its Lebesgue measure and then extending uniquely to all arcs via transverse isotopy. See [CF24b, Section 5] for more details.
@@
The fact that the length of $\lambda$ on $X$ is equal to the area of $\mathcal{O}(X,\lambda)$ combined with the $\operatorname{Mod}_{g}$-equivariance of $\mathcal{O}$ imply that it restricts to a bijection
@@
of the same name. The following omnibus theorem combines the main theorems of [Mir08] and [CF24b] together with a special case of the main result of [CF24a]. See the footnotes for specific citations.
@@
    $\mathcal{O}$ takes earthquake flow to horocycle flow in a time-preserving way: $\mathcal{O}\circ\operatorname{Eq}_{s}=u_{s}\circ\mathcal{O}$. [^4] Theorem A of [CF24b]; see also Theorem 1.1 of [Mir08].
@@
    Each $k$-pronged singularity of $\operatorname{\mathcal{O}_{\lambda}}(X)$ corresponds to an order (k2)(k-2) zero of $\mathcal{O}(X,\lambda)$ and vice versa. [^6] Proposition 5.10 of [CF24b].
@@
    If $\mathcal{Q}^{1}$ is a component of a stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$ and $q\in\mathcal{Q}^{1}$ has no horizontal saddle connections, then $q$ is a point of continuity for $\mathcal{O}^{-1}|_{\mathcal{Q}^{1}}$. [^7] A stronger version of this is proved as Theorem 15.6 of [CF24a].
@@ 3.1. Twist tori
Let $\gamma=(\gamma_{1},\ldots,\gamma_{n})$ be a multicurve on $S$; throughout the paper we use $S\setminus\gamma$ to denote the (possibly disconnected) compact surface with boundary obtained by removing an open annulus about each $\gamma_{i}$. Associated to $\gamma$ there is a cut-and-glue fibration that expresses a hyperbolic structure on $S$ in terms of the induced hyperbolic structure on $S\setminus\gamma$ and how these structures are glued together. Let $\mathcal{T}(S\setminus\gamma)$ denote the space of marked [^8] Here, the marking does not record twisting about the boundary. hyperbolic structures with totally geodesic boundary on $S\setminus\gamma$ so that for each $i$, the two boundaries of $S\setminus\gamma$ that are glued along $\gamma_{i}$ have the same length. Then we have the following principal $\mathbb{R}^{n}$-bundle:
@@ 3.2. Arc systems and ribbon graphs
[…]
@@
where $\mathcal{TRG}(S\setminus\gamma)$ is defined analogously to $\mathcal{T}(S\setminus\gamma)$. More specifically, each curve of $\gamma$ corresponds to two different cycles in the ribbon graph spines; the lengths of these cycles must match. See [AHC22, pg. 39].
[…]
@@
The space $\mathcal{TRG}(S\setminus\gamma)$ has a natural piecewise-linear structure and is stratified by the combinatorial type of the ribbon graphs. To avoid confusing these with strata of differentials, we will call such strata facets. [^10] The space of all ribbon graphs on $S\setminus\gamma$ is naturally a union of open orthants, and the gluing conditions cut out linear subspaces of each orthant. Top-dimensional facets correspond to trivalent metric ribbon graphs and lower-dimensional ones correspond to ribbon graphs with higher valence vertices — this combinatorial structure is also readily apparent in the dual picture in terms of weighted filling arc systems. For a given ribbon graph $\Gamma$ of the same topological type as $S\setminus\gamma$, we set $F(\Gamma)$ to denote the facet of $\mathcal{TRG}(S\setminus\gamma)$ consisting of metrics on $\Gamma$ satisfying the gluing conditions. Luo’s work [Luo07] implies that $\operatorname{\mathsf{Sp}}$ restricts to a real-analytic homeomorphism between each facet and a corresponding subvariety of $\mathcal{T}(S\setminus\gamma)$.
@@ 3.3. Horospheres and slices
Each horosphere has a natural lift $\mathcal{P}H_{\gamma}(\mathbf{h})$ to $\mathcal{P}^{1}\mathcal{T}_{g}$ with $\mathbf{h}\gamma$ as the second coordinate. Both the horosphere and its lift are homeomorphic to $\mathbb{R}^{6g-7}$ and are foliated by (lifts of) twist tori.
Following Forni [For21], we define a horospherical measure to be any probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ that is absolutely continuous with respect to the Lebesgue class with continuous density and such that its restriction to each earthquake flow line is the restriction of Lebesgue to an interval. We also say that the pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$ is horospherical. [^12] For example, in [AH21, Liu20, AHC22], such measures are obtained by taking certain $\operatorname{Stab}(\mathbf{h}\gamma)$-invariant measures on $\mathcal{P}H_{\gamma}(\mathbf{h})$, taking a local pushforward to the intermediate quotient $\mathcal{P}^{1}\mathcal{T}_{g}/\operatorname{Stab}(\mathbf{h}\gamma)$, then pushing down to $\mathcal{P}^{1}\mathcal{M}_{g}$.
@@
As before, there is a natural lift to $\mathcal{P}^{1}\mathcal{T}_{g}$, both the horosphere slice and its lift are homeomorphic to a Euclidean space, and both are foliated by twist tori. We say that a probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ is a horosphere slice measure if it is in the class of Lebesgue with continuous density (now on the slice) and the conditional measures on earthquake flow lines are Lebesgue on intervals. Just as above, we also use this term to denote pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$.
@@ 4.1. Twist tori and horospheres
[…]
@@
The differentials $q\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ are completely horizontally periodic, because their horizontal foliations are measure equivalent to $\mathbf{h}\gamma$, i.e., correspond to one another under the natural identification between $\mathcal{MF}_{g}$ and $\mathcal{ML}_{g}$ [Lev83]. With definitions as in [Wri15] or §5.1 below, the flat twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is just the orbit of any such by the abelian group $\mathbb{R}^{\gamma}$ of twists in the horizontal cylinders, and $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$ is Lebesgue on $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. We observe that $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is locally an affine submanifold in period coordinates, cut out by equations of the form $\operatorname{Im}(z)=h_{i}$ and $z=\ell(e)$, where $e$ is an edge of $\mathsf{Y}$ with length $\ell(e)$. By our convention that $\bm{\ell}\cdot\mathbf{h}=1$, the intersection of these affine subspaces lies in the unit area locus.
The union of the horizontal saddle connections of any $q\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is exactly $\mathsf{Y}$. Thus, the angle around each singularity of the flat metric is $\pi$ times the valence of the corresponding vertex of $\mathsf{Y}$; see also §2 of [CF24b]. However, in order to identify the ambient stratum containing $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ we also need to understand if $q$ is globally the square of an abelian differential. This can be checked combinatorially, as follows.
Definition 4.1.
Let $\Sigma$ be a compact surface with non-empty boundary and let $\Gamma$ be a ribbon graph with topological type $\Sigma$. We say that $\Gamma$ is orientable if its edges can be oriented so that each component of $\partial\Sigma$ inherits a consistent orientation. Given a multicurve $\gamma\subset S$ and a ribbon graph $\Gamma$ with topological type $S\setminus\gamma$, the pair $(\gamma,\Gamma)$ is jointly orientable if each component of $\Gamma$ can be oriented in a way that induces a consistent orientation on $\gamma$.
Clearly, if $(\gamma,\Gamma)$ is jointly orientable, then each component of $\Gamma$ is orientable. On the other hand, $(\gamma,\Gamma)$ need not be jointly orientable even if each component of $\Gamma$ can be oriented. See Figure 4.
Figure 4. Left: a pair of multicurves on a surface. Right: An orientable ribbon graph $\Gamma$ with topological type $S\setminus\gamma$ where $(\gamma,\Gamma)$ is not jointly orientable. If one builds a square-tiled surface whose vertical and horizontal cylinders are the pair of multicurves on the left, then its graph of horizontal separatrices is the one shown on the right.
Lemma 4.2.
Let $q$ be a completely horizontally periodic quadratic differential with horizontal multicurve $\gamma$ and let $\Gamma$ denote the ribbon graph consisting of its horizontal separatrices (neither the height of the curves of $\gamma$ nor the length of the edges of $\Gamma$ are relevant). Then $q$ is globally the square of an abelian differential if and only if $(\gamma,\Gamma)$ is jointly orientable.
Proof.
Consider the orientation double cover $\hat{q}\to q$. We will show that, for any component $\Sigma\subset S\setminus\gamma$, the preimage of $\widehat{\Sigma}$ in $\hat{q}$ is disconnected if and only if the component of $\Gamma$ corresponding to $\Sigma$ is orientable, and $\widehat{S}$ is disconnected if and only if $(\gamma,\Gamma)$ is jointly orientable.
By construction, each curve of $\gamma$ can be realized as a core curve of a horizontal cylinder of $q$. Cutting along these core curves results in a flat cone surface with boundary homeomorphic to $S\setminus\gamma$. This (possibly disconnected) surface now deformation retracts onto $\Gamma$ by collapsing the segments of its vertical foliation.
From this description, it follows that each component of $\Gamma$ is orientable if and only if the restriction of the horizontal foliation $\operatorname{Im}(q)$ to the corresponding component of $S\setminus\gamma$ is orientable (see also the discussion in Section 7.2 of [CF24b]), which happens if and only if the preimage $\widehat{\Sigma}$ is disconnected. Similarly, a joint orientation on $(\gamma,\Gamma)$ is equivalent to a global choice of orientation of $\operatorname{Im}(q)$ on $S$, which implies the holonomy double cover is disconnected. ∎
@@
For later use, we record how twist tori and their associated measures interact with the $P$ action.
@@
which is actually a lift of the $1$-level set of the extremal length function $L_{\mathbf{h}\gamma}:\mathcal{T}_{g}\to\mathbb{R}_{>0}$ [CF24b, GM91]. Using Theorem 2.1, $\mathcal{Q}H_{\gamma}(\mathbf{h})$ is a leaf of the unstable foliation. For any facet of $\mathcal{TRG}(S\setminus\gamma)$ we can also define the extremal length horosphere slice
By construction of $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ and Theorems 2.1 and 2.3, $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ is actually a leaf of the unstable foliation in the ambient stratum containing it. Observe that, just as in the hyperbolic setting, $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ is foliated by twist tori.
@@ 4.2. Affine invariant subvarieties
An affine invariant subvariety (or AIS) $\mathcal{L}\subset\mathcal{QM}_{g}$ is a $\mathsf{GL}^{+}_{2}\mathbb{R}-$invariant immersed sub-orbifold that is locally cut out by $\mathbb{C}$-linear equations defined over $\mathbb{R}$ in period coordinates. Its unit-area locus $\mathcal{L}^{1}$ is $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant, and comes equipped with an ergodic $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant probability measure $\nu_{\mathcal{L}}^{1}$. This measure has the property that, if $m$ is Lebesgue measure on $\mathbb{R}$, then $dm\,d\nu_{\mathcal{L}}^{1}$ is locally an affine linear measure on the defining linear subspaces in period coordinates; see [EM18, Definition 1.1]. While we will not use it directly, we mention that Filip proved that every AIS is an algebraic variety [Fil16].
@@
Let $\mathcal{Q}^{1}$ be any component of any stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$.
@@
    (Measure classification): Every $P$-invariant ergodic measure on $\mathcal{Q}^{1}$ is also $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant and is $\nu_{\mathcal{L}}^{1}$ for some AIS, of which there are at most countably many.
@@
    (Orbit closures): For any $q\in\mathcal{Q}^{1}$, the orbit closure $\overline{P\cdot q}=\overline{\operatorname{\mathsf{PSL}}_{2}\mathbb{R}\cdot q}$ is (the unit area locus of) an AIS.
@@
    (Genericity): For any $q\in\mathcal{Q}^{1}$, the orbit $P\cdot q$ equidistributes in its closure $\mathcal{L}^{1}$. That is, for any $f\in C_{c}(\mathcal{Q}^{1})$ and any $r\in\mathbb{R}\setminus\{0\}$,
    $\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\int_{0}^{r}f\left(g_{t}u_{s}(q)\right)\,ds\,dt=\int_{\mathcal{L}^{1}}f\,d\nu_{\mathcal{L}}^{1}.$
Applying Theorem 4.4 to the setting of Lemma 4.3, we deduce that the time-averaged pushforwards of a twist torus equidistribute to its PP–orbit closure.
@@
$\frac{1}{T}\int_{0}^{T}(g_{t})_{*}\left(\nu_{\gamma}(\mathsf{Y},\mathbf{h})\right)\,dt\to\nu_{\mathcal{L}}^{1}$
as $T\to\infty$, where $\nu_{\mathcal{L}}^{1}$ is the unique affine probability measure supported on $\mathcal{L}^{1}:=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})}.$
@@
$\overline{P\cdot q}=\overline{P\cdot\mathcal{Q}\mathbb{T}^{0}}=\mathcal{L}^{1}$